Moving FEROS to the MPG/ESO-2.2m Action Items Summary of Complete Items No Responsible Stat Short Summary === ==================== ==== ======================================== 001 John Pritchard ConT When is Max-Planck Time? 002 John Pritchard Comp BIAS/TCS requirements 003 John Pritchard Comp Budget access 004 Gerardo Ihle Comp Fibre Lengths 005 Gerardo Ihle Comp Room quotes 006 Alain Gilliotte Comp Vibration Measurements 007 Alain Gilliotte Comp MicroLens 008 Wolfgang Eckert Comp Rotating FEROS feasibility study 009 Oli Hainaut Comp Feasibility of TCCD for fibre head viewer 010 Ivo Saviane ConT Commissioning time justification 011 Ivo Saviane Comp Is Humidity a significant factor? 012 Alain Gilliotte Comp Final Optical Design 013 Wolfgang Eckert Comp Final Mechanical Design 014 Ivo Saviane Comp FCU GUI control panel design specification 015 Andreas Kaufer Clos Atmospheric Dispersion Corrector Design 016 Rene Mendez Clos Have Observatory priorities assigned 017 Rene Mendez Comp Telescope balance and 2.2m Pointing model 018 Luis Wendegast Comp Cooling for ACE Box + Power Supply design 019 Aldo Pizzaro Comp FEROS FibreHead Viewer TCCD 020 Jaime Alonso ConT Bypass and remove Coude interlocks 021 Carlos Ebensperger Comp Verify load bearing capacity of FCC Room floor 022 Eugenio Ureta Comp FEROS Climate Controled Room 023 Gerardo Ihle ConT Rotator gear 024 Gerardo Ihle ConT Replacing the WFI screws 025 Alain Gilliotte ConT New fibres 026 Alain Gilliotte ConT Purchase new NDF 027 Gerardo Ihle ConT Contract Climatrol for FEROS CC Room CC 028 Rodrigo Olivares ConT FWACU cables 029 Rodrigo Olivares Comp Ethernet ports and power outlets in FEROS Room 030 Gerardo Ihle Comp Moving FEROS rehersal 031 Percy Glaves ConT 2.20 WS user accounts 032 Ismo Kastinen Comp FEROS FHV-TCCD SW 033 John Pritchard Comp New BIAS-PC HDs and E-net cards 034 Percy Glaves Clos Reconfiguration of wfacu 035 Gaetano Andreoni Comp Cabling FEROS I 036 Gaetano Andreoni Comp Cabling FEROS II 037 Gerardo Ihle Comp FCU wall mount 038 Percy Glaves Comp 2p2 Operator account on wfwacu 039 Wolfgang Eckert ConT Modify Adapter Tower 040 Alain Gilliotte ConT Improve Sky Science Fibre efficiency === ==================== ==== ======================================== ConT :: Completed on Time Comp :: Completed Clos :: Closed Complete Items ============== AINumber : 001 Short Description : When is Max-Planck Time? Action Required : Check if there will be Max-Planck time during the relevant installation time. Log: 28/04/2002 jpritcha From http://www.eso.org/observing/proposals/lasilla-ins.html MPI time is December 11, 2002 --> January 10 2003. Responsible : John Pritchard Status : Completed Deadline : 30/04/2002 Deadline Status : Completed Ontime ------------------------------------------------------------------------- AINumber : 002 Short Description : BIAS/TCS requirements Action Required : Provide the requirements of the TCS interface to the BIAS CCD image acquition software. Log: 07/08/02 jpritcha BIAS/TCS requirements, along with the overall proposed computing plan for the project were presented to F.Gutierrez in a meeting 07/04/02. 08/27/02 jpritcha Ismo has written a VLT-standard interface to allow BIAS to retrieve the TCS status from w2p2tcs. Full testing will only be possible during October commissioning. 09/26/02 jpritcha This item is complete. Responsible : John Pritchard Status : Closed Deadline : 26/10/2002 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 003 Short Description : Budget access Action Required : Contact Otmar Stahl to find out the exact details of the budget and the procedures for access to the funds. Log: 30/04/2002 isaviane Here is the O.S. answer to John's message: "Dear John, Thank you for the mail. It is good to hear that Feros will get a new life at the 2.2m telescope. Currently we still have exactly 40 210,04 Euros left. I do not know exactly what you need, but this money should be enough. I already had an e-mail exchange with Jorge Melnick about the procedure to follow for using the Feros budget for moving Feros to the 2.2m telescope. The only constraint I have, is that our administration will make a transaction only when they have an invoice. This means that you, unfortunately, cannot directly access the account, which I would prefer. What we could do, and what I proposed to Jorge, is, that you (ESO, La Silla or Garching) order what you need and send me the invoice. I will then ask our administration to pay the bill. If you have another proposal - I am open for anything as long as it is ok with the administration on all sides." 06/04/02 jpritcha As discussed at the progress meeting today, whenever possible the procedure will all preliminary work will be done by the person involved up to and including obtaining Quote and Proforma. This will then be provided to the Project Manager (Pritchard) who will forward the Proforma to Otmar Stahl. He will then place the order specifying the delivery address as ESO Warehouse Garching (who will forward it to La Silla), and make payment on the Invoice. Email sent to O.Stahl specifying this procedure and asking if he accepts it. 06/05/02 jpritcha Here is the exact procedure accepted by Otmar Stahl: 1) The La Silla person involved collects all relevant "preliminary" documentation and finally obtains a Proforma. 2) The Proforma is provided to Project Manager for checking/acceptance and then is forwarded by the Project Manager it to Otmar Stahl, usually by FAX. 3) Based on the Proforma Otmar Stahl places the order, specifying the delivery address as per the Proforma (which in most cases will be the ESO-Garching Warehouse. Of course the invoice address will be Otmar Stahl, Landessternwarte Hidelberg-Konigstuhl. 4) Upon reciept of Invoice Landessternwarte Hidelberg-Konigstuhl pays it, forwarding a copy if the invoice to the Project Manager. 5) Of course any other documentation Otmar Stahl and/or Landessternwarte Hidelberg-Konigstuhl receive wrt an order should also be copied to the Project Manager. The FAX number for Otmar is +49-6221-509-202. Responsible : John Pritchard Status : Completed Deadline : 30/04/2002 Deadline Status : Closed ------------------------------------------------------------------------- AINumber : 004 Short Description : Fibre Lengths Action Required : Measure the required Fibre Length for the three location options. Log: 29/04/2002 jpritcha In a meeting today between Pritchard, Ihle, Eckert & Saviane we estimated the distance required by the 'internal' route just to get from the fibre head to the exit point of the telescope pier was ~12m. Assuming then that we are using the current fibres it is then not possible to take this route for neither the Pier nor the TIO options (it was never considered an option for the dome floor option). Nor is it a possibility even if we have new fibres but a restricted to a length of not more than 15m (as asserted by Kaufer). Therefore the implication is clear. The internal route for the fibre is not possible. 06/04/02 jpritcha As discussed at the progress meeting today, according to AGi "exact" measurements (using a "rope") are to be made during the next day or two by Ernesto Araya. Only the TIO room option is required as at this date this is the only location option still being considered. 06/27/02 agilliot Fiber length has been measured. The distance from the Cassegrain to the Telescope fork structure to the base and to the Feros new room location gives 16m with subtancial length reserved for the telescope moves and at the instrument entrance. If the exact present length is 13m then we can have a problem !!! but the solution will be to drill an hole from the telescope floor to directly the Feros room .... Then we can gain the Fork/base length ...... Responsible : Gerardo Ihle Status : Completed Deadline : 02/05/2002 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 005 Short Description : Room quotes Action Required : Obtain quotes for environmentally controlled rooms for each of the three locations. For each location require a quote for the following levels of climate control: Temperature Humidity ----------- -------- 16+/-0.5 <65% 16+/-0.1 <65% 16+/-0.5 ~30+/-5% 16+/-0.1 ~30+/-5% Log: 29/04/2002 jpritcha In a meeting today between Pritchard, Ihle, Eckert & Saviane, Ihle asserted that he expects to complete by the end of this week. Prelimianry discussions with Climatrol indicate control to with +/-5% at 20-30% humidity will not be feasible, however 50+/-5% may be feasible. Another option is "Variation smoothing" rather than strict control. This may be an option if 50+/-5% proves beyond reason. 06/17/02 jpritcha At the Design Review meeting held 2002-Jun-11 Gerardo Ihle presented the following preliminary quote from Climatrol based on requirements suggested at the PDR: Temperatura ambiente interior: 18,0 +/- 0,3 oC(*) Humedad relativa interior: 45 +/- 5 %(*) Disipacion de calor interior constante de maximo 500 W a) Suministros nacionales, instalacion, programacion y puesta en marcha: UF 380 b) Suministro de controles y humidificador: Valor Ex Fabrica (Suiza): Euro 3.200 (mas transporte e internacion) Based on the Design Review discussions the requirements for the FCCR have been revised as follows: T = 16 +/- 0.5 C 30% < H < 60% with the understanding that H will not be strictly controlled as such, but merely monitored and extra humidty will be added to the FCCR whenever the H falls below 30%. Would also be useful to know how cost goes with delta-T. If the difference between +/-0.5 and +/-0.3 is not significant, why not??? 07/01/02 jpritcha G.Ihle reports: Subject: RE: FEROS Action Item 005 Date: Thu, 27 Jun 2002 17:49:48 -0000 From: Gerardo Ihle To: "'John Pritchard'" John, I just received an e-mail of Climatrol. The person to study the case is back from a trip so he will be looking to the alternative asap. Best regards, Gerardo 07/06/02 jpritcha Further news from G.Ilhe, Climatrol have provided a quote as follows: Temperature: 17.0 +/- 2.0 C Variation of T: +/- 0.5 C/h Cost: UF 247.80 ~= 16300*247.80 = CLP 4,039,140 ~= CLP/672.829 = EUR 6003 G.Ihle wrote: > Herewith you will find the Climatrol quotation. Regarding the variation of > +/-0.5degC represents that same one that was proposed in the original > project for the installation in the 1.52m. This of course is better in > reality. The second figure that I found strange was the 20w dissipation > inside the FEROS room, but they told that as in the other figure, it came > from the original quotation. This is not important if we will have the same > equipment inside this room. Given that Climatrol are the company currently responsible for the AC system at the ESO-1.52m which implies that they know the capabilities of that equipment well, proceeding with this quote would seem to be the safest option. G.Ihle has been asked to obtain detailed plans for the system proposed by Climatrol. 07/30/02 jpritcha Superceded by AI#027. Responsible : Gerardo Ihle Status : Completed Deadline : 10/05/2002 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 006 Short Description : Vibration Measurements Action Required : Measure the vibrations due to dome movement for the three location options plus at the current FEROS room at the ESO-1.52 as a reference. Log: 06/04/02 jpritcha Report on measurements made at 1.5m and 2.2m telescopes submitted by AGi 23 May 2002. Report however recommends new measurements to be made in TIO's room on floor surface without carpet. Also recommends the piles directly into the rock solution. 06/04/02 jpritcha As discussed at the progress meeting today, the carpet was today removed, therefore AGi should be able to make new vibration measurements within one or two days. 06/27/02 jpritcha At the DR-11/06 the new vibration measurements by AGi were discussed implying the proposed location is good. Awaiting final report. 06/27/02 agilliot The second alternative place for the Feros room is ok. Report is ready to be issued (small pb on my computer due to the file size!) 07/08/02 jpritcha Final report delived by A.Gilliotte in email dated Sun, 30 Jun 2002 11:12:41 -0400. Responsible : Alain Gilliotte Status : Assigned Deadline : 09/05/2002 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 007 Short Description : MicroLens Action Required : Make optical design and place order for a rod microlens with a 2.3" fibrehead aperture. Log: 06/04/02 jpritcha As discussed at the progress meeting today, the rod microlens is no longer an option as neither supplier suggested by A.Kaufer can commit to manufacturing the required lens. Therefore we revert back to a ball lens design. This AI is therefore closed. Responsible : Alain Gilliotte Status : Closed Deadline : 09/05/2002 Deadline Status : Closed ------------------------------------------------------------------------- AINumber : 008 Short Description : Rotating FEROS feasibility study Action Required : Feasibility study for mounting the FEROS adapter on the rotator. Log: 29/04/2002 jpritcha In a meeting today between Pritchard, Ihle, Eckert & Saviane, Eckert indicated a physical inspection of the 2.2m adapter would be a good idea to insure there are no 'design surprises' not indicated or even not included in current WFI plans, however we all agreed it was probably best not to do this during MPI time. The current MPI allocation ends noon May 19th and is followed by one VA run and then Idle time 22/05-->29/05. It should thus be done before the final design (which will be dur end of May). We also noted that the WFI Grism has the potential to conflict with the WFI/FEROS selector mirror assembly. The solution is that FEROS simply will not be usable while the grism is mounted. Design must be checked to ensure that the WFI/FEROS selector mirror assembly will not clash with grism even whilst in rest position. 06/04/02 jpritcha As discussed at the progress meeting today, rotating FEROS is now considered integral to the project, not merely an option. Therefore it must be incoproated into the final mechanical design. This AI can therefore be closed. Responsible : Wolfgang Eckert Status : Closed Deadline : 09/05/2002 Deadline Status : Complete ------------------------------------------------------------------------- AINumber : 009 Short Description : Feasibility of TCCD for fibre head viewer Action Required : Check the availability/feasibility of using one of the NTT TCCDs as the fibre head viewer. Check also for the availability/feasibility of the ACE box (NTT) and LCU (old EMMI). Log: 06/04/02 jpritcha According to emails from PSi and OHa (both dated 05/07/02) we have: *) One TCCD (R.Parra) *) The ACE Box (P.Sinclaire) *) An incomplete LCU, including VME crate, CPU, Sequencer and I/O cards ( the old EMMI Red LCU) (P.Sinclaire). Therefore still required: *) ACE-head and LCU-ACE cables *) TCCD Fibre Interface and 24V Power Supply. 06/04/02 jpritcha As discussed at the progress meeting today, the ACE-head cable and the LCU-ACE fibre are now in preparation by Rolando Medina, although he awaits more info regarding required lengths (amongst other things??). We resolved that the ACE box would be mounted on the telescope but NOT on the rotator. Therefore the length required for the ACE-head cable is AT LEAST 2.5m. Jaime Alonso assures the good performance of TCCDs with a cable of this length. The LCU-ACE fibre on the other hand must reach from the ACE mounted on the telescope to the LCU which is forseen to be installed in the TIO room (henceforth to be known as the 2.2m FEROS room, or FEROS room when it is clear that we are talking about the 2.2m FEROS room and not the 1.52m FEROS room.) Since the LCU will be in the FEROS room, we will therefore require the "Large" VME crate. The TCCD itself is of the "small format" variety (approx 12mmx8mm ==> 2.34'x1.56'). One implication of this is that this CCD can NOT be readout in a window format, therefore this TCCD will NOT be suitable for guiding, only for field acquisition and verification of object on fibre. The above mentioned 24V Power Supply has yet to be sourced, though J.Alonso is checking to see if he has a "spare one" somewhere. This power supply will need to be mounted on the telescope along with the ACE box. Cooling for the ACE box and/or Power Supply must be considered. An AI (for L.Wendegast has been created). The already existing WFI VME cooling supply can be tapped into. The above mentioned TCCD Fibre Interface was in the end supplied by the NTT. Software for the TCCD needs to be provided by SWC, but should be "trivial" as it should require only "standard" VLT TCCD drivers and software, including RTD. Clearly at this point the feasibility of the exNTT-TCCD is well established and this AI can therefore be closed. A new AI will therefore be opened for implementation of the FEROS FibreHead Viewer TCCD (FFV-TCCD). Responsible : Oli Hainaut Status : Closed Deadline : 09/05/2002 Deadline Status : Complete ------------------------------------------------------------------------- AINumber : 010 Short Description : Commissioning time justification Action Required : Write a plan to justify the request of commissioning time. Log: 30/04/2002 isaviane plan sent to ohainaut@eso.org Responsible : Ivo Saviane Status : Completed Deadline : 30/04/2002 Deadline Status : Completed on time ------------------------------------------------------------------------- AINumber : 011 Short Description : Is Humidity a significant factor? Action Required : Determine if and at what level variations of humidity affect FEROS stability. Suggest to contact FEROS observers who may be able to comment. And to perform a simple test -- Make a initialisation, and then take ThAr's as the humidity in the FEROS room is artificially raised (not above 65%). The humidifier in the 152 control room could be used for this purpose. Log: 04/28/02 jpritcha Suggested observers who may be able help: Jens Viggo Clausen (who is the one observer as far as I know who has raised the issue of humidity control) and Luca Pasquini and Licio Da Silva who have been/are FEROS BIG programme PIs. 06/04/02 jpritcha On 04/30/02 ISa contacted Jens Viggo Clausen (hereafter JVC) via email. 06/04/02 jpritcha In an email dated 09 May 2002 ISa wrote: Dear all: I did some tests of how humidity can affect measurements with FEROS. First of all, a bit of "theoretical" work. The first 4 figures show the dependence of the refractive index of air vs. humidity and temperature, at wavelength = 300 nm and 900 nm. The calculations vs. T have been made at 30% humidity, those vs. H have been made at 30 C. The index is computed at an altitude of 2km, assuming a scale height of the atmosphere of 7km. The figures show that - n obviously depends on lambda - the variations of n do not depend very much on lambda - the dependence of n on T is ca. twice that on H, so - since Herman Hensberge found a cycle of amplitude 600m/s and period of 30min, corresponding to the cycle in temperature of amplitude 1.5 K, - we expect a change of the same size for a change of 3% in humidity (but changes in humidity do not cycle on such short timescales) Just to have an idea of the effect. If a beam of light enters the Feros room with an inclination of 45deg, a change of 100% in humidity will deviate it by ca. 6 microns after a path of 10 meters. --- Then the experimental work. I put the control room humidifier inside the FEROS room, and took HeAr arcs while the humidity was increasing. The last 2 figures show the displacement of a line along the dispersion. I averaged the line along crossdispersion, and then for each arc I inserted its line into a new image, in a different position. So the rows in the figure are pixels along the wavelength and the columns are the number of the arc, where rows up to 4 were taken at constant humidity ca. 14%, and then the humidity increases up to 44% at row 12. One can see that a change of 30% in the humidity shifts the lines by roughly 1/2 pixel along the dispersion (i.e. 7 microns). 06/04/02 jpritcha In an email dated 10 May 2002 JVC replied to ISa writing that he had not done any significant studies of the influence on changing humidity on his science data. ISa then forwarded to JVC the above report. 07/08/02 jpritcha The fact that FEROS is "Not a high precision radial velocity machine" (DR-11/06) means that the possibility to control strictly the humidity is simply beyond the scope of this project. Responsible : Ivo Saviane Status : Assigned Deadline : 31/05/2002 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 012 Short Description : Final Optical Design Action Required : Finalise Optical Design Log: 06/04/02 jpritcha Note, ADC requires a 5th position for the calibration unit sliding mirror. 06/04/02 jpritcha As discussed at the progress meeting today, with the rod microlens option not possible we revert back to a ball lens design. 1mm diameter, saphire, off the shelf ball lenses of high quality (according to ???) are available for just 10Euro each. Furthermore the ball lens allows a larger diameter aperture in the fibre head, up to 2.2arcsec, implying no ADC required for airmasses up to ~1.8 (AGi estimate). Therefore we adopt the following: *) For Period-70 we manufacture a fibre head with 2.2arcsec aperture. *) If and when in the future an ADC is implemented, part of that project should be a new fibre head with a smaller entrance aperture. As for the final optical design itself, including ball lens option, AGi assures it is ready. Final acceptance is reserved for late June when A.Kaufer will hopefully be able to make a final appraisal. 06/27/02 jpritcha At the DR-11/06 we reverted back to a 1.8arcsec fibre entrance aperture, the overiding consideration being maintaning the present Resolving Power (R) of FEROS. 07/08/02 jpritcha Note that an ADC thus makes no difference to the desiable entrance aperture, in order to maintain R=48000, the entrance aperture must always be 1.8" at the 2.20m. 09/27/02 jpritcha Design phase completed by end of July. Responsible : Alain Gilliotte Status : Completed Deadline : 31/05/2002 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 013 Short Description : Final Mechanical Design Action Required : Finalise Mechanical Design Log: 06/04/02 jpritcha As discussed at the progress meeting today, the Final Mechanical Design is well underway, with all components being at least ~60% finalised. Furthermore in order to meet the final deadline, fabrication of some components (in particular the arm) is already underway in the workshop. The deadline for delivery of the mechnical components is Mid-August which will allow 2 weeks for integration, and one month for integration of the associated electronics. The electronics is being handled by J.Alonso while the control software for FEROS is being done by R.Olivares. The two moving components for the FEROS adapter (the mirror arm and the calibration sliding mirror) will be controlled by the same VME as will be used for the FCU. The mirror arm will have only two functions (in and out) and its status will be detected by limit switches. No position encoder is forseen. The arm will be operatable in both local (i.e. manual) and remote (i.e. computer controlled) modes. Torque must constantly be applied to hold the mirror in position even in the worst case of gravity acting directly along the axis of motion. Note, as of July 1st, R.Olivares moves to the SEST, and is available for FEROS work only for 50% of his time. The following days and nights of technical/reserved time are available at the ESO-1.52m for testing of the control of the FCU and the 2.2m-FEROS adapter: 27/06 (1n), 29-30/08 (2n). There is also 1 half night of reserved time available on 09/08 if it is useful for anything. 31/08 --> 28/09 is Brazillian time. The last night of observations at the ESO-1.52m (a half night with FEROS) is 29/09 (other half of night is reserved time). Note, examination of the 2.2m and the dismounted WFI revealed only one unexpected aspect. Two interlocks probably from the "Coude days" need to be bypassed so that they can be removed. An AI for J.Alonso has been created. 06/27/02 jpritcha At the DR-11/06 WEc reported that the design between 60 and 80% complete for the different subcomponents. A further 100hrs being required to complete the design, 200hrs for preparation of detailed drawings and 500-600 ours for fabrication. 07/01/02 jpritcha W.Echert reports: Subject: FEROS status Date: Thu, 27 Jun 2002 09:17:53 -0400 From: "Wolfgang Eckert" To: "John Pritchard" The final designs of following units are ready: 1.) Mirror arm with drive 2.) Mirror housing 3.) TCCD unit 4.) Small mirror carriage (90%) 5.) Fiber head unit with calibration optics 6.) Fiber head (90%) 6.) Big Rotator Flange 7.) Main optics support Detail drawings of following units are ready and given to the workshop: 1.) Mirror arm with drive 2.) Mirror housing 3.) TCCD unit 4.) Small mirror carriage (0%done) 5.) Fiber head unit with calibration optics 6.) Fiber head (80%) 7.) Big Rotator Flange 8.) Main optics support Following units are in the workshop: 1.) Mirror arm with drive (ready) 2.) Mirror housing (ready) 3.) TCCD unit (under work) 4.) Fiber head unit with calibration optics (0%done) 5.) Big Rotator Flange (10%done) 6.) Main optics support (0%done) The mirror housing must be slightly modified because of the new size of the mirror. Delivery time of the new mirror: 6 weeks. Following parts or units have to be designed: 1.) Boring pattern for the big flange 2.) Modification of the calibration unit 3.) Fixation of the calibration unit 4.) ND filter wheel unit 5.) ACE box fixation All materials are ordered and the half of them arrived already. The rails for the 'small mirror carriage' have also a delivery time of 6 weeks. 09/27/02 jpritcha Design phase completed by end of July. Responsible : Wolfgang Eckert Status : Completed Deadline : 31/05/2002 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 014 Short Description : FCU GUI control panel design specification Action Required : Provide Design specification to ROl Log: 06/16/02 jpritcha Deadline set as per email from Ivo: Subject: Re: New Action Item Date: Fri, 14 Jun 2002 01:06:46 -0400 From: Ivo Saviane To: John Pritchard John Pritchard wrote: >Then maybe we should allow at least two tournos... so I guess this will mean >some time in July, so I will make the deadline 31/07, OK? According to the Q3 schedule, my first turno will be on Aug 9... Responsible : Ivo Saviane Status : Completed Deadline : 09/08/2002 Deadline Status : Pending ------------------------------------------------------------------------- AINumber : 015 Short Description : Atmospheric Dispersion Corrector Design Action Required : Design an Atmospheric Dispersion Corrector Log: 06/04/02 jpritcha In an email dated 28 Apr 2002, AKa pointed out the desirability of an Atmospheric Dispersion Corrector given the then design limitations of the optical design. Andreas wrote: After thinking about the design a bit I find that the basic design is pretty easy (no raytracing needed ...): the telescope f/8 beam has to be converted into a f/4.6 beam for the fibers. That's all. with a fiber core of 90mu this gives: 90mu * 8.0 / 4.6 = 156.5mu aperture on sky 156mu * 0.0117"/mu = 1.83" aperture on sky If we could get perfect centering of the lens, 100mu of the fiber could be used, corresponding to 2.0" on sky. The only way to increase the aperture on sky is to increase the fibre diameter which will result in the fact that the image slicer could not be used anymore and the resolving power would be reduced. This seems not acceptable to me. 1.8" corresponds to the atmospheric dispersion between 370-900nm at airmass 1.5. (at the 1.52 the numbers were 2.7" corresponding to airmass 2). To limit observations to AM<1.5 seems to me quite a limitation. So the only way out of this dilemma is to think about a simple, small and efficient atmospheric dispersion corrector. If you want, I can have a look into this in Santiago. Here I really need to use the raytracing program to optimize over the wavelength range. I will also have to check the rod lens design because something must be wrong with the numbers which ended up in Allain's report. 06/04/02 jpritcha In an email dated 13 May 2002 ISa suggested ADC be planned for in current opto-mechanical design, but not implemented for Nov-2002, but rather for the beginning of (say) ESO Period 71 (i.e. April 2003). This suggestion is endorsed by RMe & WEc. 06/04/02 jpritcha Note, ADC requires a 5th position for the calibration unit sliding mirror. 06/04/02 jpritcha As the ADC is nolonger considered a part of the "Moving Project" this AI can be closed. An ADC and a new FibreHead will be considered as a seperate future project. Responsible : Andreas Kaufer Status : Closed Deadline : 30/05/02 Deadline Status : Closed ------------------------------------------------------------------------- AINumber : 016 Short Description : Have Observatory priorities assigned Action Required : Ask management to clearly define observatory priorities, particularly for the workshop Log: 06/04/02 jpritcha In an email date 11 May 2002, RMe asked JMe & OHa to assign priorities. 06/27/02 jpritcha In a conversation Rene confirmed that "management" had confirmed the top priority status for Moving FEROS. Responsible : Rene Mendez Status : Completed Deadline : 30/05/02 Deadline Status : Closed ------------------------------------------------------------------------- AINumber : 017 Short Description : Telescope balance and 2.2m Pointing model Action Required : Once the new FEROS optics and mechanics are installed on the 2.2m, the telescope will need to be balance and the pointing model will need to be recomputed. Log: 06/14/02 jpritcha In an email Rene wrote: Subject: Re: FEROS Action Item #017 Date: Fri, 14 Jun 2002 12:50:51 -0400 From: "Rene A. Mendez" To: John Pritchard No problem , John. 10/31/02 jpritcha Telescope balances have been established for both the case of FEROS/WFI adapter only AND FEROS/WFI adapter + WFI. For FEROS the current pointing model is Ok -- during Commissioning-III (16/10/2002 --> 25/10/2002) very bright objects object were just off the field of view of the FFHV TCCD but easy to locate. Responsible : Rene Mendez Status : Completed Deadline : 10/31/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 018 Short Description : Cooling for ACE Box + Power Supply design Action Required : Design cooling system for ACE Box + Power Supply Log: 06/04/02 jpritcha As discussed at the progress meeting today, the already existing WFI VME cooling supply can be tapped into. 06/11/02 jpritcha Cooling of course needs to be provided for the TCCD (not the ACE Box + Power Supply). One point especially to bear in mid is that the water pipes to tthe TCCD must be insulated. 07/06/02 jpritcha In email dated Fri, 05 Jul 2002 12:05:35 -0500, L.Wendegast reports: +++ After a meeting with P.Sinclaire we conclude that there is not problem to make this work. Moreover,there are two alternative to feed water to these units. One is to have a similar system than in the other units (a Neslab Chiller, that we do not have available) or a direct feeding by connecting Fiera and these units in serie, using water-glycol solution. The last alternative is recommended by the hardware maintenace manual of these units. +++ Clearly the 2nd option should be used. L.Wendegast requested to draw up detailed plans. 07/16/02 jpritcha In email dated Wed, 10 Jul 2002 11:43:50 -0500, L.Wendegast provided a detailed plan. Responsible : Luis Wendegast Status : Completed Deadline : 07/12/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 019 Short Description : FEROS FibreHead Viewer TCCD Action Required : Implement the FEROS FibreHead Viewer TCCD Log: 06/04/02 jpritcha In coordination with SWC the complete FEROS FibreHead Viewer TCCD system (FFV-TCCD) should be implemented asap. The LCU including VME crate (Large), CPU, Sequencer and I/O cards (including the TCCD Fibre Interface) and the ACE box have been sourced (mainly from the NTT). As discussed at the progress meeting today, the ACE-head cable and the LCU-ACE fibre are now in preparation by Rolando Medina, although he awaits more info regarding required lengths (amongst other things??). We resolved that the ACE box would be mounted on the telescope but NOT on the rotator. Therefore the length required for the ACE-head cable is AT LEAST 2.5m. Jaime Alonso assures the good performance of TCCDs with a cable of this length. The LCU-ACE fibre on the other hand must reach from the ACE mounted on the telescope to the LCU which is forseen to be installed in the TIO room (henceforth to be known as the 2.2m FEROS room, or FEROS room when it is clear that we are talking about the 2.2m FEROS room and not the 1.52m FEROS room.) Since the LCU will be in the FEROS room, we will therefore require the "Large" VME crate. This will probably need to be at least the length determined by the "rope measurements" to be made by Ernesto Araya (see AI-004). The TCCD itself is of the "small format" variety (approx 12mmx8mm ==> 2.34'x1.56'). One implication of this is that this CCD can NOT be readout in a window format, therefore this TCCD will NOT be suitable for guiding, only for field acquisition and verification of object on fibre. A 24V Power Supply for the ACE box has yet to be sourced, though J.Alonso is checking to see if he has a "spare one" somewhere. This power supply will need to be mounted on the telescope along with the ACE box. Cooling for the ACE box and/or Power Supply must be considered. An AI (for L.Wendegast has been created). The already existing WFI VME cooling supply can be tapped into. Software for the TCCD needs to be provided by SWC, but should be "trivial" as it should require only "standard" VLT TCCD drivers and software, including RTD. The exact configuration of the observing environment and therefore which workstation the TCCD should be controlled from has yet to be decided. 06/11/02 jpritcha Subject: Re: FEROS Action Item #019 Date: Tue, 11 Jun 2002 18:00:18 -0400 From: Aldo Pizarro To: John Pritchard Hi John, I agree with the responsibility and with the deadline cheers Aldo 06/11/02 jpritcha In an email Aldo wrote: Subject: Re: FEROS Action Item #019 Date: Tue, 11 Jun 2002 18:00:18 -0400 From: Aldo Pizarro To: John Pritchard Hi John, I agree with the responsibility and with the deadline cheers Aldo 07/01/02 jpritcha A.Pizarro reports: Subject: Re: FEROS Action Item #019 Date: Thu, 27 Jun 2002 10:57:11 -0400 (SAT) From: Aldo Pizarro To: j.pritchard@eso.org I have talked with Ismo and, as a first step, we will make some test with the Tech. CCD software during the week end of the next turno. I already have the Tech CCD VME implemented and tested offline. Respect to the maximun lenght of the cables from head I could not find an official documentation but according with the experience of Paranal, the maximum lenght that they are using is 3 meters so, 2.5 meters it is good. any other quetion, please go ahead. 09/27/02 jpritcha Peter Sinclaire reports that the hardware is all ready and awaits only a final in situ test to check for unforseen problems. 10/31/02 jpritcha System installed an successfully tested during Commissioning-III. Responsible : A.Pizarro Status : Completed Deadline : 07/31/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 020 Short Description : Bypass and remove Coude interlocks Action Required : Bypass and remove Coude interlocks Log: 06/04/02 jpritcha Next available idle time at 2.2m is 3 days starting 21 June but according to R.Mendez this is "a bit tight". After that the next available idle time is 7 nights starting 17 Sep. 06/05/02 jpritcha In an email Jaime wrote: Subject: Re: FEROS Action Item #020 Date: Wed, 05 Jun 2002 08:24:55 -0400 From: Jaime Alonso To: John Pritchard no problem, but from June 18 till 25 I am in PAO. I guess the best are the 7 days starting on September 17. 09/26/02 jpritcha Completed successfully during the Spetember 2.2m idle time. Responsible : Jaime Alonso Status : Closed Deadline : 09/23/02 Deadline Status : Completed on Time ------------------------------------------------------------------------- AINumber : 021 Short Description : Verify load bearing capacity of FCC Room floor Action Required : Verify that the FEROS Climate Control Room floor can bear the load of the FEROS instrument. Log: 06/17/02 jpritcha The FEROS instrument weighs in total approximately 900kg. Therefore we should verify that 250kg per leg can be withstood by the floor. Log: 07/06/02 jpritcha In an email dated Thu, 04 Jul 2002 17:42:19 -0400 C.Ebensperger wrote: > I have verified the floor of the room of Feros and he has an > approximate resistance of 190 to 210 Kg/cm2 is studying the > possibility to incorporate a special product in the floor in order to > diminishing the vibrations, improve the resistance and to even at the > same time. Good news. The measured resistence should be perfectly adequate to accomodate FEROS. However NO "special product" should be applied to the floor. FEROS should sit directly on the concrete. 07/08/02 jpritcha A final check of the area of each leg in contact with the floor must be made to ensure that it is greater than ~1.5cm2, but it is hard to imagine it is less than this! Responsible : Carlos Ebensperger Status : Completed Deadline : 07/10/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 022 Short Description : FEROS Climate Controled Room Action Required : Design and build the FEROS Climate Controled Room Log: 07/03/02 jpritcha Prepare detailed plans for the FEROS Climate Controled Room -- an enclosure with dimensions of l x w x h = 2.5mx3.4mx2.0m (height is a lower limit), constructed of polystyrene insulated steel panels of 75mm thickness (the same type as used for refrigerated chambers). Please see Section 3 of the document "Moving FEROS from the ESO 1.52m Telescoep to the MPG/ESO 2.2m Telescope" for further details. http://www.ls.eso.org/lasilla/Telescopes/2p2T/E2p2M/FEROS/FEROSto2.2/ 07/22/02 jpritcha At the FDR-18/07 C.Ebensperger reported that the quote for the FEROS Climate Control Room (excluding climate control) was due during the week of 22/07-28/07. Once contracted the process should take 2 weeks for preparation (by supplier) and then 1 week for assembly (on site). Despite this, due to other committments the earliest completion date possible is mid-September. 09/26/02 jpritcha Responsible : Eugenio Ureta Status : Closed Deadline : 07/12/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 023 Short Description : Rotator gear Action Required : Remove a small section from the rotator gear to allow the passage of cable Log: 07/16/02 jpritcha In an email dated Sun, 14 Jul 2002 09:16:09 -0400 W.Eckert wrote: > I accept this AI. The work will be done, when we install the first time the > rotator flange. Therefore the Deadline is set accordingly. 07/20/02 jpritcha At the FDR-18/07 it was realised this must be done during the September 2.2m idle time (17th --> 23rd). Therefore since W.Eckert will be on vacation this AI must be reassigned to G.Ihle. 09/26/02 jpritcha Completed successfully during the Spetember 2.2m idle time. Responsible : G.Ihle Status : Closed Deadline : 09/23/02 Deadline Status : Completed on Time ------------------------------------------------------------------------- AINumber : 024 Short Description : Replacing WFI screws Action Required : Replacing WFI screws Log: 07/08/02 jpritcha With the extra distance between the mounting plate and WFI introduced by the FEROS/WFI adapter, the current bolts fixing WFI to the mounting plate will be too short. They must therefore be replaced by longer bolts. 07/16/02 jpritcha In an email dated Wed, 10 Jul 2002 07:52:55 -0000, G.Ihle wrote: > As I'm going in holidays and just coming back at the endo f the month > I will appreciate if discuss this point with Wolfgang. 07/20/02 jpritcha At the FDR-18/07 it was realised this must be done during the September 2.2m idle time (17th --> 23rd). Therefore since W.Eckert will be on vacation this AI must be reassigned to G.Ihle. 09/26/02 jpritcha Completed successfully during the Spetember 2.2m idle time. Responsible : G.Ihle Status : Closed Deadline : 09/23/02 Deadline Status : Completed on Time ------------------------------------------------------------------------- AINumber : 025 Short Description : New fibres Action Required : Investigate and if possible pruchase and prepare new FibreHead-FEROS fibres Log: 07/22/02 jpritcha At FDR-18/07 it was realised that routing the FibreHead-FEROS fibres directly through a "hole in the floor" (necessary if the current 13-14m fibres are used) presents too great a safety risk, in particular in relation to safety of the fibres themselves. Therefore we must secure new FibreHead-FEROS fibres. These MUST be obtained in time that they can be prepared and still be ready for October 18th, when they will be installed on the telescope. 07/30/02 jpritcha In a an email from A.Gilliotte: From: "Alain Gilliotte" To: "John Pritchard" Subject: Re: New FEROS Action Item #025 Date: Wed, 24 Jul 2002 18:54:13 -0400 John, PR issued for the science (100m) and calibrations fibers - Polymicro type so action item closed Responsible : Alain Gilliotte Status : Completed on Time Deadline : 10/18/02 Deadline Status : Completed on Time ------------------------------------------------------------------------- AINumber : 026 Short Description : Purchase new NDF Action Required : Purchase new NDF Log: 07/22/02 jpritcha Current NDF has optical density range 0-2 (100-1% transmission). Some thought should be given to whether the new one should be 0-2 or 0-3. 07/30/02 jpritcha In a an email from A.Gilliotte: From: "Alain Gilliotte" To: "John Pritchard" Subject: Re: New FEROS Action Item #026 Date: Wed, 24 Jul 2002 18:55:04 -0400 John, PR issued for a ND wheel from Coherent, 152.4mm diameter and density range 0.0-2.0 action item closed Responsible : Alain Gilliotte Status : Completed on Time Deadline : 09/15/02 Deadline Status : Completed on Time ------------------------------------------------------------------------- AINumber : 027 Short Description : Contract Climatrol for FEROS CC Room CC Action Required : Act on quote from Climatrol Log: 07/22/02 jpritcha During FDR-18/07, it was realised that the quote from Climatrol for the the climate control of the FEROS Climate Controled Room must be acted on with all possible due haste. The following schedule must be aimed for: 15/09: Structure of FEROS CC Room will be ready. 30/09: Components currently installed at the ESO-1.52 become available for relocation to MPG/ESO-2.20m 06/10: Installation of climate control must be complete. Note: **) The climate control installation should be complete as soon after 30/09 as possible, therefore whereever possible components should be installed prior to 30/09 so that the minimum amount of work remains to be done after 30/09. **) Components not dependent on those components currently installed at the ESO-1.52 nor on the structure of the FEROS CC Room should be installed asap. **) Components not dependent on those components currently installed at the ESO-1.52 but dependant on the structure of the FEROS CC Room should be installed as soon as the FEROS CC Room is ready, which according to the above, should be 15/09. 07/30/02 jpritcha In an email from G.Andreoni: Date: Thu, 25 Jul 2002 10:40:24 -0400 From: Gaetano Andreoni To: John Pritchard CC: Gerardo Ihle Subject: Re: New FEROS Action Item #027 Dear John, I made a first evaluation of the quotation. The good news are that it is not that much work and one month advance notice is the minimum required. That is, the Order can be handled by Gerardo at the beginning of August. Also, it seems to me that the best place for the Fan-Coil is in the telescope basement room, hanging from the ceiling. This will ensure the shortest ducts and pipes, and no vibration transmitted to the Feros foundation. The bad news are that there is no air conditioning in the outer Feros room. The equipment installed there are for ventilation only. Cheers Gaetano So back to Gerardo... 08/27/02 jpritcha In arecent email G.Ihle wrote... Subject: Traslado de FEROS a nueva sala del 2.2m Date: Wed, 14 Aug 2002 11:51:57 -0000 From: Gerardo Ihle To: "Clemente Chappuzeau (E-mail)" Clemente, Debo comunicarte algunos cambios que han sucedido con el proyecto: 1) La nueva sala estara terminada por ISG el 15 de Septiembre. 2) Entre el 15 y el 30 de Septiembre se pueden realizar trabajos en la sala que no impliquen equipo en uso en el 1.52m 3) El equipo de climatizacion de FEROS estara en uso en el 1.52m HASTA el 30 de Septiembre. 4) El equipo debera estar instalado (listo) en la nueva sala del 2.2m el 6 de Octubre. 5) Entre el 6 y el 9 de Octubre, el equipo debera estar funcionando de tal manera de limpiar la sala de posible polvo en suspension y para permitir una calibracion de este. 6) El 9 de Octubre, el instrumento FEROS es trasladado a la nueva sala en el 2.2 Te mando esta informacion para que planifiques tu venida a La Silla en una nueva fecha ya que la sala recien sera entregada a Eugenio el 26 de Agosto. De ahi tienen que iniciar su instalacion en el 2.2m Saludos, Gerardo 09/26/02 jpritcha The Climatrol guys are here now working on the system, a sure indication that this AI was successfully completed! Responsible : G.Ihle Status : Closed Deadline : 10/06/02 Deadline Status : Completed on Time ------------------------------------------------------------------------- AINumber : 028 Short Description : FWACU cables Action Required : Design FWACU cables for fabrication by LED Log: 07/22/02 jpritcha Bear in mind that according to the measurements made by A.Gilliotte "The distance from the Cassegrain to the Telescope fork structure to the base and to the Feros new room location gives 16m with subtancial length reserved for the telescope moves and at the instrument entrance." Therefore it is probably a good idea to specify a length of 20m for all cables needing to go from FEROS Room to adapter. 07/30/02 jpritcha In an email from R.Olivares: Date: Wed, 24 Jul 2002 15:18:25 -0400 (SAT) To: rmedina@eso.org From: Rodrigo Olivares Subject: Cables proyecto FEROS Cc: rolivare@eso.org, jpritcha@eso.org Hola Rolando, Respecto a los cables sobre los que conversamos el mes pasado y que necesitamos para el proyecto "FEROS at the ESO220", te envio el detalle (todos son cables estandard MACCON a Motor/Tacho/Encoder como los usados en Harps): Funcion Largo Tipo ========================================================= Cable1 NDFW 5 mts circular (NO limits sw) Cable2 Shutter 5 mts circular (NO limits sw) Cable3 SCSM 20 mts lineal (2 limit switches) Ademas te recuerdo obtener las especificaciones de la funcion que esta haciendo Jaime (FWSM) con un largo de 20 metros. El plazo limite para el que los necesito es el 31 de Agosto, siendo el mas prioritario el Cable 3 (SCSM). Saludos, Rodrigo The original task of this AI is therefore complete, but it should be extended to include follow-up with LED to ensure the cables are delivered ontime. Deadline will therefore be appropriately extended. 10/07/02 jpritcha Rodrigo wrote: Subject: Re: [Fwd: FEROS Action Item #28] Date: Thu, 3 Oct 2002 14:17:06 -0700 (Pacific Daylight Time) From: "Rodrigo Olivares" To: All the cables are ready. Responsible : R.Olivares Status : Completed Deadline : 08/31/02 Deadline Status : Completed on Time ------------------------------------------------------------------------- AINumber : 029 Short Description : Ethernet ports and power outlets in FEROS Room Action Required : Define requirements, request work, followup Log: 07/30/02 jpritcha This AI is already underway: Date: Wed, 24 Jul 2002 16:45:08 -0400 (SAT) To: John Pritchard From: Rodrigo Olivares Subject: Re: 2.2m ethernet ports Hi, I asked for 7 ports in the FCU room: 4 at the wall where the FCU will be installed (VME chassis) and 3 more in the front wall (WS and BIAS PC). Moreover, I asked to maintain the 2 ports in the corner of the FEROS camera. I hope it is clear for you now. If you want I can explain you in more detail. Saludos, Rodrigo. PD: In a similar way I asked for the Power (220V) connections. 10/07/02 jpritcha Inspected today, work appears to be complete. I assume it was thoroughly yet tested by the technicians who installed it. Responsible : R.Olivares Status : Completed Deadline : 08/31/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 030 Short Description : Moving FEROS rehersal Action Required : Reherse moving the FEROS table Log: 08/14/02 jpritcha ISG has constructed from wood a frame representing the FEROS table top to allow a rehersal moving the table top from 1.52 to 2.20. As discussed with Gerardo today, he will work out the details of this during is tourno in the first part of September. 10/31/02 jpritcha Rehersal made and plan formulated. Moreover the plan was successfully implemented with FEROS being successfully transported from 1.52 to 2.20 with a minimum of fuss within a period of ~4hrs! Responsible : Gerardo Ihle Status : Completed Deadline : 09/30/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 031 Short Description : 2.20 WS user accounts Action Required : Create feros & ferosmgr accounts on 2p2 WSs Log: 08/14/02 jpritcha Create feros & ferosmgr user accounts on the 2.20m Workstations as per the Version 0.4 of the plan document. 08/17/02 jpritcha Completed as per email... Subject: Re: New FEROS Action Item #031 Date: Thu, 15 Aug 2002 09:58:49 -0400 From: Percy Glaves To: John Pritchard John, The accounts were created at w2p2tcs w2p2oh & w2p2off the password for all of them is ... OK I'll tell you that by phone! On w2p2drl were also created ... but 1 problem, group 50 in drl is used by ftp. So they got their private groups. If this is a problem please contact Marcus. Responsible : Percy Glaves Status : Completed Deadline : 15/08/02 Deadline Status : Completed on Time ------------------------------------------------------------------------- AINumber : 032 Short Description : FEROS FHV-TCCD SW Action Required : Install, configure etc software for the FEROS Fibre Head Viewer Technical CCD Log: 08/14/02 jpritcha As discussed with J.Ibsen & A.Pizarro today, the FHV-TCCD is ready for installation at the 2.20. Software installation at the 2.20 should be done during the September Idle time (Sep 17-24th noon-noon). This is the only available idle time between now and the start of commissioning which would obviously be too late! Note that the FEROS/WFI Adapter will NOT be ready for the camera head to be mounted therefore the SW will have to be installed whilst the camera is on a testbed. 08/19/02 jpritcha Discussions with I.Kastinen and F.Gutierrez on 08/17 derived the following plan: **) The Fibre Head Viewing Technical CCD will be physically moved to the 2.2m asap, and definitely no later than Sep 16th. **) I.Kastinen will begin work on preparing the Fibre Head Viewing Technical CCD software directly on the 2.2m TCS WS during daytime as soon as the hardware is at the 2.2m telescope. **) The full Fibre Head Viewing system (hardware and software) WILL be ready for commissioning to start on October 16th. No special presence of I.Kastenen is required, except during the FEROS Commissioning (October 16-21). A change of tourno of I.Kastinen to suit requirements will therefore be required. 09/27/02 jpritcha Ismo wrote: Subject: Re: FEROS Action Item #32 Date: Fri, 27 Sep 2002 10:50:03 -0400 From: Ismo Kastinen To: John Pritchard John Pritchard wrote: > Can you please give me a quick update on progress on the FEROS Fibre Head Viewer > Technical CCD Software Item. > 1) I will start generating and testing TCCD software tomorrow, if team gives me day time of telescope. 2) The new (J5600) TCS WS failed during last idle time, so the C200 was changed back. The offsets sent to telescope were not executed. I will try to find out the reason in this turno. What I heard from my team mates other functions of TCS were fine. That's all. 10/02/02 jpritcha Ismo wrote: Subject: FEROS Action Item #32 status Date: Thu, 03 Oct 2002 12:37:30 -0400 From: Ismo Kastinen To: John Pritchard CC: SWC Team here is detailed report of FEROS Fibre Head Viewer Technical CCD Software status. All the following is installed (if applicable) only to J5600 WS. The C200 does not have any changes. - 2.2m TCS Feros SW integration module has been created. I will use pkgin tool to generate Feros software. John: I need to know names and versions of software modules which I should include there. This module is not yet archived due I am waiting Garching to create it. - ESO/MPI 2.2m TCS Feros SW Regeneration document first version was written. Document is not yet archived due I am waiting Garching to create it. - both WS and LCU environment for camera were created. Unfortunately, I could not test them in the telescope due I could not connect to LCU in ODT lab. I don't know reason for this. Even ping was not answering from telescope. However, in the wlsops2 were created environments to test camera (thanks Ruben and Peter) and camera is working fine. - the offset problem detected last idle time with new WS (J5600) was tested. I suppose, I found solution. The new WS is sending commands for offset to VME too fast, so VME does not have enough time to change to correct state to be able to execute offsets. I did not have time to perform thorough tests, but adding delays to WS process sending offset commands seemed to help. No offset failed after that. This has to be tested well during commissioning. - there was an additional problem during last idle time with differential guiding which I heard just wednesday. The speeds entered to telescope control panel was not corresponding values which differential guiding was actually performed. I could not test this problem, due information was given so late to me. However, we have lot of time to correct this behavior. Fernando Selman said to me that next run with differential guiding will be couple of months from now. As a backup solution we will keep C200 workstation as a spare in the telescope until we can solve this problem. 10/31/02 jpritcha System successfully installed and tested during Commissioning-III. Responsible : Ismo Kastinen Status : Completed Deadline : 10/16/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 033 Short Description : New BIAS-PC HDs and E-net cards Action Required : Purchase 20Gb HDs and 100Mb E-net cards for the two BIAS PCs Log: 09/26/02 jpritcha Following email was sent to Flavio Gutierrez: Subject: FEROS @ 2.2m :: New hardware for BIAS PCs Date: Tue, 27 Aug 2002 22:02:09 -0400 From: John Pritchard To: "Gutierrez-Willer, Flavio" Flavio, As we discussed during my previous tourno, some new hardware is needed for the BIAS PCs for FEROS @ the 2.2m: 2 x 20Gb IDE HardDisks. 3 x 100Mb PCI ethernet cards. The above provides for one of each for each of the two BIAS PCs, plus one spare incase of failure (we already have one 20Gb HD in the BIAS PC in use at the ESO-1.52m). I would appreciate it very much if you would arrange for purchase of these items from your "usual supplier". Please charge the account 210002C of work package 644 (Medium Sized Telescopes Mainteneance - manager Rene Mendez). Cheers John No response as yet... 10/31/02 jpritcha Hardware delivered by Gutierrez. Responsible : John Pritchard Status : Completed Deadline : 09/30/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 034 Short Description : Reconfiguration of wfacu Action Required : Reconfigure wfbc into wfwacu Log: 09/26/02 jpritcha The C110 workstation currently installed at the ESO-1.52m telescope as wfbc must be reconfigured as wfwacu. A complete fresh installation is forseen with HPUX 11.00 + VLTSW-MAR2002-CCSLite. The existing harddisks should be backed up before doing anything else. The WS will be available from Oct 05 until Oct 06. After installation it should be delivered to the 2.20m telescope for installation in the FEROS Room. Please coordinate with Rodrigo Olivares at the time. Installation must include creation of the feros and ferosmgr user accounts in addition to the standard VLT accounts applicable for that machine. 10/31/02 jpritcha The dependance of the B&C Spectrograph CCD VME for booting from wfbc made it too risky to contemplate reconfiguration at this time. Instead we will reconfigure w1p5ins --> wfwacu once w1p5ins becomes free from its duties at the ESO-1.52m at the end of the year. That will then be a normal La Silla Action Item. Responsible : Percy Glaves Status : Closed Deadline : 10/06/02 Deadline Status : Closed ------------------------------------------------------------------------- AINumber : 035 Short Description : Cabling FEROS I Action Required : FEROS Climate Control Room cabling Log: 09/26/02 jpritcha Two phases of cabling are forseen to be necessary, the first (this AI) should be done on approximately Oct 9th or 10th (in coordination with the Reassembly and Alignment of FEROS, contact John Pritchard) and involves cables and fibres enterring the FEROS Climate Control Room. The second (AI#36) should be done on approximately Oct 16th (in cordination with installation of the FEROS/WFI adapter, contact John Pritchard) and involves cables running from telescope to the FEROS Room and the finalisation of the FEROS Climate Control Room cabling. The following "cables" (along with their dimensions) are forseen to enter the FEROS Climate Control Room from the FEROS Room: 1) Liquid Nitrogen auto feed 2) CCD controller to BIAS PC fibre optic cables 3) FEROS environmental monitoring sensor cables 4) Adapter to FEROS Science fibre optic cables It is forseen that 1)-3) can be done on a permanent basis (ie including light and airtight sealing) on the Oct 9. 4) should be prepared but not finalised until the 2nd phase. Between Oct 9 and Oct 16 4) will be used to feed the FEROS calibration unit fibres directly to FEROS. It owuld therefore be great, but not essential (read "if it is easy to do then go ahead but otherwise don't worry about it) if temporary light and airtightness cound be provided. All preparartions (ie drilling holes in the walls of the FEROS Climate Control Room) should preferably be completed BEFORE Oct 6. 09/28/02 jpritcha Sorry I forgot to include the cable dimensions above. After inspecting the 1.52m FEROS room today I can provide the following probable specifications for the required holes. 1) Liquid Nitrogen auto feed: approximately 7cm diameter hole. The size of this hole at the 1.52m should be measured accurately. Also the height from the ground should be the same as currently at the 1.52m. This can be checked from the room immediately before the actual 1.52m FEROS room. The hole should be made in the Western wall (i.e. the wall with the door) near the North Western corner of the room. The hole should be as far from the corner as possible without interferring with the movement of the door. 2) Cables associated with the liquid Nitrogen auto feed system: approximately 4cm diameter. This hole should be 10cm directly above the Liquid Nitrogen auto feed hole. 3) Science Fibre cables: Assuming a casing similar to the one currently in use then a hole of approximately 2cm diameter will be required. This hole should probably be located in the ceiling of the room along the EAST-WEST center line approximately 0.6m from the western wall. 4) All other cables: A hole of approxiamtely 5cm diameter. This hole should also be located on the Western wall 5cm down from the ceiling and 5cm from the NW corner. In case that any of the descriptions above are not clear or for any other reason, you prefer to wait for me to return on Oct 6th before proceeding on this. 10/02/02 jpritcha G.Andreoni wrote: Subject: Re: FEROS Action Item #035 :: Update Date: Tue, 01 Oct 2002 09:15:22 -0400 From: Gaetano Andreoni To: John Pritchard CC: Gerardo Ihle Making these holes is quite straightforward and I would do them at the time of the installation of the corresponding pipe/cable. For me the best timing would be after installing the table (so we know the exact location of everything) and before reinstalling the optical elements. By the way, do you have an action for the fixation of the four legs to the floor? It's a heavy job and preparation must be completed before moving the table. ...and... Subject: [Fwd: FEROS Action Item #035 :: Update] Date: Tue, 01 Oct 2002 15:36:37 -0400 From: Gaetano Andreoni To: John Pritchard CC: Gerardo Ihle After talking with ODT people, they can provide the position for the holes (continuous flow and fibers, 1. & 2.) now. The mounting will be as devised by me, that is after the table and before the optical elements. 31/10/02 jpritcha Completed. Responsible : Gaetano Andreoni Status : Completed Deadline : 10/10/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 036 Short Description : Cabling FEROS II Action Required : Telescope to FEROS Room cabling Log: 09/26/02 jpritcha Two phases of cabling are forseen to be necessary, the first (AI#35) should be done on approximately Oct 9th or 10th (in coordination with the Reassembly and Alignment of FEROS, contact John Pritchard) and involves cables and fibres enterring the FEROS Climate Control Room. The second (this AI) should be done on approximately Oct 16th (in cordination with installation of the FEROS/WFI adapter, contact John Pritchard) and involves cables running from telescope to the FEROS Room and the finalisation of the FEROS Climate Control Room cabling. The following "cables" are forseen to run from telescope to FEROS room and/or FEROS Climate Controled Room: 1) FEROS Technical CCD cable 2) FEROS Science fibre optic cables 3) LCU to Adapter cable 4) FEROS Calibration Unit fibre optic cables Cables should be routed from adapter to Fork mount (via the Dec axle), along the fork to the pier, down the south face of the pier, through the floor and into the FEROS Room. A loop must be provided between fork and pier to allow the full motion of the telescope. 31/10/02 jpritcha Completed. Responsible : Gaetano Andreoni Status : Completed Deadline : 10/18/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 037 Short Description : FCU wall mount Action Required : Design and construct a simple wall mounting for the FEROS Calibration Unit Log: 09/26/02 jpritcha Design and construct a simple wall mounting for the FEROS Calibration Unit in the FEROS Room. 31/10/02 jpritcha A floor standing support was contructed and installed during Commissioning-III. Responsible : Gerardo Ihle Status : Completed Deadline : 10/09/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 038 Short Description : 2p2 Operator account on wfwacu Action Required : Create standard e2p2ops account on wfwacu Log: 11/01/02 jpritcha Please create an e2p2ops account on wfwacu which is needed for the creation/maintenance of FEROS configuration files. 12/22/02 jpritcha Completed. Responsible : P.Glaves Status : Completed Deadline : 11/13/02 Deadline Status : Completed ------------------------------------------------------------------------- AINumber : 039 Short Description : Modify Adapter Tower Action Required : Design and implement modifications of adapter tower Log: 11/01/02 jpritcha In light of the results of Commissioning-III, some `slight' modifications are required to the FEROS/WFI Adapter Tower. These modifications will incorporate the fact that the Fibre Head needs to be ~3cm further from the M3 than in the original plan plus fixing the distance between TCCD and TCCD-lens to ease the process of focusing. Alain Gilliotte will provide the corrected optical design. It is expected that correct alignment of the optics, in particular insuring that the calibration fibres are incident on the fibre head at the correct angle will greatly improve throughput in the blue for the calibration lamps. 12/22/02 jpritcha Modifications were completed in time for Commissioning-IV. In essence, the tower was cut in two lengthened by inserting a spacer block. Responsible : W.Eckert Status : Completed on Time Deadline : 11/13/02 Deadline Status : Completed on Time ------------------------------------------------------------------------- AINumber : 040 Short Description : Improve Science Fibre efficiency Action Required : Align apertures, ball lens and fibres Log: 11/01/02 jpritcha During Commissioning-III it was found that the fibre efficiency maximum acheivable fibre efficiency was only ~8% where as at the ESO-1.52m ~15% was typically achieved in excellent conditions. Furthermore the relative fibre efficiency was found to be ~1.8, whereas at the 1.52m it had been ~1.2. It is belevied that this is due to poor fibre head aperture, ball lense and fibre alignment. 12/22/02 jpritcha Significant improvements were made and the efficiency was raised to 11% by fabricating new fibre ferules and ball lense holders. The problem now appears to be the alignment and size (220um instead of 158um) of the fibre head apertures. A new FibreHead will need to be fabricated (AI#41). Responsible : A.Gilliotte Status : Completed on Time Deadline : 11/13/02 Deadline Status : Completed on Time